OpenStudy (anonymous):

in a triangle ABC, AB =sqrt2, AC = 1/sqrt2 and the angle at A is 60 deg, Find the length of BC and the size of the angle at C.

OpenStudy (anonymous):

use cosine law.

OpenStudy (anonymous):

any chance you could show how

OpenStudy (baldymcgee6):

http://www.youtube.com/watch?v=pGaDcOMdw48

OpenStudy (campbell_st):

BC uses the Cosine Rule or laws of Cosine... the Angle C is found using the Sine Rule or Laws of Sine start by drawing a diagram  1st find BC using the Cosine Rule remember a = BC c = AB and b = AC \[a^2 = b^2 + c^2 - 2ab \cos(A)\] substitute and evaluate. 2. Find angle C using the Sine Rule \[\frac{\sin(C)}{c} = \frac{\sin(A)}{a}\] again substitute and evaluate